You have calculated the following returns for an individual stock:

Periods               Return

Month 1                 12%

Month 2                   8%

Month 3                 9%

Month 4                 10%

What is the stock's expected return and standard deviation?

Expected return: (round your answer to the fourth decimal, e.g. if the answer is 1.11%, you should input: 0.0111)

Standard Deviation:   (round your answer to the fourth decimal, e.g. if the answer is 1.11%, you should input: 0.0111)
Answer #1

We can calculate the expected return using the following formula:-

Expected return (Er)= \sum return/no of periods

= (12+8+9+10)/4

= 9.75%

= 0.0975

Standard deviation can be calculated by taking route of the following

= Sum of {(X-Er)2/ no of period}

Month return (X) X-(Er) X(Er)2

1

12 2.25 5.0625
2 8 -1.75 3.0625
3 9 -0.75 0.5625

4

10

0.25

Total

0.0625

8.75

So standard deviation = 8..75/4 = 2.1875

Route of 2.1875 =1.479%

ie, 0.01479

..... Hope you are satisfied with the solution, kindly support by rating the answers.......